1 + sin theta = 2 cos^2 theta (solve on the interval [0,2pi]
Answer by lwsshak3(11628)

1 + sin theta = 2 cos^2 theta (solve on the interval [0,2pi]
1+sinx=2cos^2x
1+sinx=2(1-sin2x)
1+sinx=2-2sin^2x
2sin^2x+sinx-1=0
(2sinx-1)(sinx+1)=0
..
2sinx-1=0
sinx=1/2
x=π/6, 5π/6
or
sinx+1=0
sinx=-1
x=3π/2